### v4

@@ Abstract
Transformer neural networks (TNN) have shown great performance in natural language processing (NLP) [1], machine translation [2], computer vision [3], and other fields in recent years. While recurrent neural network (RNN) [4] and long short-term memory (LSTM) [5] models run sequential computation tasks during both training and inference, transformer facilitates high levels of computation parallelism throughout both processes using an attention mechanism. Thus, TNN is becoming a potential alternative to CNN, RNN, and LSTM [6, 7]. There are many transformer models, such as full transformers containing both encoder and decoder [8], BERT [9, 10], ALBERT [11], structBERT [12], and others. These models contain different numbers of encoder and decoder stack [8] for different applications. A single encoder will often require a latency on the order of 100s of [13]. Around 38% to 64% of this time is spent in the multihead attention (MHA) mechanism depending on the number of tokens in the input sequence [14, 15], and the rest of the time is spent on feed forward network (FFN). Unfortunately, general-purpose platforms like GPUs and CPUs often suffer from low computational efficiency, underutilized memory bandwidth, and substantial compilation overheads for MHA layers [16]. MHA and FFN disabledisabletodo: disableFFN is defined and only use once, I would trim done the number of acronyms in the paper. also occupy most of the on chip storage units [17, 18, 19]. disabledisabletodo: disablewhat does this mean? are you talking about on-chip memory? (Ehsan’s Ans.: Fixed) Therefore, it is essential to prioritize efficient hardware deployment on resource-constrained devices. FPGAs have gained widespread use for accelerating DNNs due to their high level of parallelism, high energy efficiency, and low latency disabledisabletodo: disableI would delete ”short development cycle”. I think you meant this as compared to designing an ASIC, but no many people are designing ASICs for specific neural networks anyway, (Ehsan’s Ans.: deleted) [20, 21]. Recently, some works have successfully built FPGA based custom hardware accelerators for transformers [22, 13, 18]. Application-specific integrated circuits (ASIC)-based accelerators also exist [23]. Most of these works compress the model by using different weight pruning strategies to accelerate computations and, as a result, must incorporate specialized hardware to support sparse tensor access. However, the benefit of supporting sparse tensors does not offset the added cost of this hardware disabledisabletodo: disableI modified the preceding sentence–confirm that you agree.(Ehsan’s Ans.: Confirmed).
@@
Lu et al. [22] accelerated the attention mechanism and feedforward network separately, but did not implement the full transformer encoder. Ye et al. [24] focused on accelerating only the attention mechanism using a reconfigurable systolic array for the transformer. Similarly, Zhang et al. [25] concentrated on accelerating the attention layer through hardware-software co-design. disabledisabletodo: disableIn this section you say ”some others papers” etc a lot and its hard to follow. Why not call a few papers out by name, like ”Ye et al. showed …” That would be easier to read. You also have a lot of ”transition words” like ”moreover, therefore, however”, to me, this means the section needs more structure and anyway to do that would be to call out individual papers form authors by name. [24, 25, 22] neglecting feedforward networks and non-linear operations. Some other works accelerate the full transformer networks but their logic circuits go through the time-consuming synthesis steps for different models or they perform poorly on the same model with different configurations [26]. They lack generality to support other variants.
@@
As transformer variants continue to evolve with differing parameters, designing a generic and efficient accelerator that can be customized to the structural characteristics of these variants becomes increasingly valuable. Thus, a versatile accelerator is needed that can efficiently handle dense matrix computations across various TNN applications. Digital signal processing (DSP) resources are capable of high-speed computation at higher frequencies. Proper utilization of them depends on the implementation method. For example, most accelerators [13, 27, 28, 29] used high-level synthesis (HLS) tools, while some used hardware description language (HDL) [30, 31, 32] for design. While HLS requires less implementation time compared to HDL, writing efficient HLS code to use parallel DSPs for optimal performance is challenging [33]. Additional challenges include storing the vast number of TNN parameters in the on-chip memories of FPGAs, which typically have a size of 5MB for low-end devices such as the ZCU104 and 35MB for high-end devices such as the Alveo U200 [34].
A challenge is executing the extensive number of multiplication and accumulation (MAC) operations required by TNNs on the DSPs, with Ultrascale+ FPGAs offering approximately 9024 DSPs. Therefore, input matrices must be partitioned into tiles. However, developing an optimal partitioning scheme that aligns well with the architecture presents a significant challenge. The data access and computation patterns differ across various blocks within the transformer, which also prevents acceleration. Therefore, assigning a dedicated hardware module to each block allows easier design and optimization.
In this paper, we utilized HLS tool to design a programmable accelerator for transformers. Our HLS code is optimized to maximize the parallel usage of DSPs. The computations are handled by DSP48 slices, with the required data stored in BRAMs, LUTRAMs, or registers within the processing elements (PE). Our architecture features efficient tiling for both the attention mechanism and linear transformations, enabling enhanced parallel computations and communication to maximize the acceleration of transformers.
The paper contributes in the following ways to the development of high-performance FPGA-based accelerators for transformer neural networks (TNNs), with an emphasis on improving computational efficiency, resource utilization, and adaptability across different hardware configurations:
@@
A novel accelerator architecture for a complete transformer that maximizes DSP and LUT utilization to enhance parallel processing and achieve low latency. disabledisabletodo: disableThis is the first time latency is discussed, maybe we should add a paragraph above to help sell the idea?
@@
An analytical model to estimate resource utilization and latency.
A modular design approach to account for different computation and data access patterns in different components of the TNN.
A parameterized HLS code that enables design-time adjustments, making it easier to modify the design.
A runtime adaptive feature enabling dynamic adjustment of parameters from software, enabling execution of different models without any hardware re-synthesis.
The full source code 111https://github.com/Kabir-Ehsan/Transformer_on_FPGA to reproduce the presented results or improve the design.
@@ 3.1 Load Weights Unit
There are three Load_weight units. One unit loads weights to the weight BRAMs (Fig. 2) of each attention head according to demand from external memories. Another unit loads weights to the weight BRAMs (Fig. 3) of the feedforward network. The third one loads weights for the layer normalization module. Weight BRAMs of the attention module are declared as two-dimensional arrays of size () in HLS. is tile_size for the attention module, which represents the dimension of the sub-matrices into which the larger weight matrices are divided. They are synthesized as dual-port BRAMs or LUTRAMs. Therefore, they are loaded with partial data from external memories for each tile at each iteration. Weight BRAMs of the feedforward network are declared as two-dimensional arrays of size (. TSFFN is tile_size for the FFN module which equals to . (4 TSFFN) equals to . Thus, they are tiled along two dimensions (row and column of the matrix) and loaded iteratively for each tile in each dimension. They are also synthesized as dual-port BRAMs or LUTRAMs. The weight BRAMs for the layer normalization unit are declared as one-dimensional arrays of size in HLS. They are not tiled, and so, loaded at once, and they are converted into dual-port BRAMs or LUTRAMs after synthesis. Algorithm 1 describes the weight loading process.
@@ 3.2 Load Inputs Unit
Inputs are loaded into an input BRAM from external memory. It is a dual port BRAM implemented as a two-dimensional array of size () in HLS, where is sequencelength. It is reused to store the outputs of each encoder/decoder layer so that it can send inputs to the subsequent layer. One Load_inputs unit loads data to the intermediate Input BRAMs (Fig. 2) of each attention head from external memory using algorithm 2. These BRAMs are declared as a two-dimensional array of size (). Therefore, tiling is applied along the column of the matrix, and they are replenished with new data () times. Another Load_inputs unit loads data to the Score BRAMs (Fig. 3) of the FFN1 module which are declared as a two-dimensional array of size () using algorithm 3. The third Load_inputs unit loads data to the LN1 Buffers (Fig. 3) of the FFN2 module which are declared as a two-dimensional array of size ()) using algorithm 4. Different units are used because of different numbers of computations, tile sizes, and array shapes.
@@ 3.3 Load Biases Unit
Biases are stored in registers because of their small size. There are three Load_bias units. One unit loads biases for attention head computations to the registers of each attention head from external memory described by Algorithm 5. The same algorithm 6 describes loading of biases for feedforward networks and layer normalization unit. Biases are declared as one-dimensional arrays in HLS and complete array partition pragma converts them into registers. Since they are not tiled, they are loaded once with all the data.
@@ 3.4 Activation Unit
The activation functions of the transformer are defined here. ReLU, GeLU and softmax activation functions are usually used in transformers. They are defined according to their equations and are implemented with LUTs after synthesis. The codes for ReLU and GeLU are simple, so, only the code for softmax is shown in Algorithm 7.
@@ 3.5 Layer Normalization Unit
The mean and variance of the outputs from the attention layer and feedforward network layer are calculated in Layer Normalization (LN) unit according to Equation 4 which is implemented by algorithm 8 in HLS. The values are normalized using the mean and variance, and the results are multiplied by weights and then added with biases element-wise. The outputs from FFN1 and FFN3 go through the LN unit before further processing (Fig. 3).
3.6 Attention Module
The architecture to accelerate the attention module is shown in Fig. 2. There are three main processing modules (PM) in it. They are denoted as , and according to the output they produce. The number of instances for these modules is equal to the number of attention heads (h). Each module contains an array of processing elements (PE). A PE is comprised of a DSP48 that performs multiplication and accumulation (MAC) operations. The PE array’s data access pattern and computational requirements differ across modules. Therefore, they are defined separately as functions in HLS to produce distinct sets of PE arrays. This approach enables optimization of each RTL module separately. Input data and weights are stored in multiple BRAMs/LUTRAMs to enable parallel access.
In our architecture, each PE is independent, with its own local memory, control, and computing unit. The weights (, , ) for generating query (Q), key (K) and value (V) matrix are declared as separate two-dimensional arrays of size () in HLS. The number of heads and tiles, and the array partition directive on HLS determine how the arrays will be partitioned to generate multiple two-port BRAMs. Due to the limited ports of BRAMs, array partitioning and data loading are efficiently managed to ensure that data required simultaneously by a DSP are stored in separate BRAMs. The , , and matrices of size () are stored in intermediate buffers.
3.6.1 QKVPM module:
module generates the query, key, and value matrices. This module contains the , , BRAMs, and input () BRAMs from which data is accessed in parallel by parallel DSP units. The arrays used in this module are divided into subarrays using our tiling technique to fit into the BRAMs or LUTRAMs. The number of times the module is accessed depends on the tile size. There are a total of () tiles or iterations. At each iteration, the , , , and buffers are loaded with distinct data. Then the computations start in the PEs. Simultaneously, the biases for the , , and matrices are loaded to registers from off-chip memory while the module performs computations. Then they are added with , , matrices. Algorithm 9 outlines the computations of this module where the 2nd loop (line 6) is pipelined causing the innermost loop (line 8) to be fully unrolled. This generates () PEs.
3.6.2 QKPM module:
module performs the matrix-matrix multiplication operations between the and matrices. Because these matrices are relatively small, they are not tiled. Algorithm 11 describes these operations. The innermost loop (line 6) is fully unrolled, generating () PEs for this module. This module contains the and buffers from which data is retrieved by the DSP units. As the division operation described in Equation 1 is executed within this module using LUTs, the number of parallel operations is constrained to prevent excessive use of LUTs. A matrix () of attention weights is generated within this module, which is stored in BRAM or registers. Subsequently, these values are forwarded to the non-linear softmax activation function. This function, as described in HLS, performs its computations using LUTs and FFs.
3.6.3 SVPM module:
The output matrix () derived from the softmax operation is transmitted to the module (Algorithm 12), where it undergoes matrix-matrix multiplication operations with the value (V) matrix. Algorithm 12 fully unrolls the innermost loop (line 6) , resulting in PEs. The output from this module is referred to as the attention score.
3.7 Feedforward Network Module
The architecture to accelerate the feedforward network module is shown in Fig. 3. There are three RTL modules for FFN to perform the operations of feedforward networks of different architectures. The definitions of the functions representing the modules have different dimensions of arrays for the inputs and outputs in HLS. These arrays are converted into BRAMs/LUTRAMs after synthesis. The number of computations inside each module is different, so each FFN has a separate function. They contain a different number of processing elements after synthesis because of different unrolling factors of the innermost loop. The weights are stored in a two-dimensional array () with dimensions () in HLS, where represents the tile size in the FFN. Both and are followed by layer normalization (LN) modules.
3.7.1 FFN1PM module:
module performs first linear transformation on the attention scores. The arrays used by the PEs are tiled along both dimensions. Thus, this module is accessed times to complete the operation. Algorithm 14 outlines the computations of this module, where the 2nd loop is pipelined causing the innermost loop (line 7) to be fully unrolled. This generates PEs which equals to .
3.7.2 FFN2PM module:
module performs second linear transformation on the normalized outputs of module. The arrays used by the PEs are tiled along both dimensions. Thus, this module is accessed times to complete the operation. Algorithm 15 outlines the computations of this module, where the 2nd loop is pipelined causing the innermost loop (line 7) to be fully unrolled. This generates PEs that equals .
3.7.3 FFN3PM module:
module performs final linear transformation on the normalized outputs of module. The arrays used by the PEs are tiled along both dimensions. Thus, this module is accessed times to complete the operation. Algorithm 10 outlines the computations of this module, where the 2nd loop is pipelined causing the innermost loop (line 7) to be fully unrolled. This generates PEs that equals .

### v5

@@ Abstract
Transformer neural networks (TNN) have shown great performance in natural language processing (NLP) [1], machine translation [2], computer vision [3], and other fields in recent years. While recurrent neural network (RNN) [4] and long short-term memory (LSTM) [5] models run sequential computation tasks during both training and inference, transformer facilitates high levels of computation parallelism throughout both processes using an attention mechanism. Thus, TNN is becoming a potential alternative to CNN, RNN, and LSTM [6, 7]. There are many transformer models, such as full transformers containing both encoder and decoder [8], BERT [9, 10], ALBERT [11], structBERT [12], and others. These models contain different numbers of encoder and decoder stack [8] for different applications. A single encoder will often require a latency on the order of 100s of [13]. Around 38% to 64% of this time is spent in the multihead attention (MHA) mechanism depending on the number of tokens in the input sequence [14, 15], and the rest of the time is spent on feed forward network (FFN). Unfortunately, general-purpose platforms like GPUs and CPUs often suffer from low computational efficiency, underutilized memory bandwidth, and substantial compilation overheads for MHA layers [16]. MHA and FFN disabledisabletodo: disableFFN is defined and only use once, I would trim done the number of acronyms in the paper. also occupy most of the on chip storage units [17, 18, 19]. disabledisabletodo: disablewhat does this mean? are you talking about on-chip memory? (Ehsan’s Ans.: Fixed) Therefore, it is essential to prioritize efficient hardware deployment on resource-constrained devices. FPGAs have gained widespread use for accelerating DNNs due to their high level of parallelism, high energy efficiency, and low latency [20, 21]. Recently, some works have successfully built FPGA based custom hardware accelerators for transformers [22, 13, 18]. Application-specific integrated circuits (ASIC)-based accelerators also exist [23].
@@
Lu et al. [22] accelerated the attention mechanism and feedforward network separately, but did not implement the full transformer encoder. Ye et al. [24] focused on accelerating only the attention mechanism using a reconfigurable systolic array for the transformer. Similarly, Zhang et al. [25] concentrated on accelerating the attention layer through hardware-software co-design. In contrast, ADAPTOR is developed to support the entire transformer neural network (TNN). Some other works accelerate the full transformer networks but their logic circuits go through the time-consuming synthesis steps for different models or they perform poorly on the same model with different configurations [26]. These approaches lack the generality to support diverse variants, whereas ADAPTOR eliminates the need for repeated synthesis across models.
@@
As transformer variants continue to evolve with differing parameters, designing a generic and efficient accelerator that can be customized to the structural characteristics of these variants becomes increasingly valuable. Thus, a versatile accelerator is needed to efficiently handle dense matrix computations across various TNN applications, and ADAPTOR is designed to fulfill this role. Digital signal processing (DSP) resources are capable of high-speed computation at higher frequencies. Proper utilization of them depends on the implementation method. For example, most accelerators [13, 27, 28, 29] used high-level synthesis (HLS) tools, while some used hardware description language (HDL) [30, 31, 32] for design. While HLS requires less implementation time compared to HDL, writing efficient HLS code to use parallel DSPs for optimal performance is challenging [33]. To address this, ADAPTOR employs optimized HLS coding techniques. Additional challenges include storing the vast number of TNN parameters in the on-chip memories of FPGAs, which typically have a size of 5MB for low-end devices such as the ZCU104 and 35MB for high-end devices such as the Alveo U200 [34] and executing the extensive number of multiplication and accumulation (MAC) operations required by TNNs on the DSPs, with Ultrascale+ FPGAs offering approximately 9024 DSPs. Therefore, input matrices must be partitioned into tiles. However, developing an optimal partitioning scheme that aligns well with the architecture presents a significant challenge, one that has been carefully addressed in the design of ADAPTOR. The data access and computation patterns differ across various blocks within the transformer, which also prevents acceleration. To overcome this, ADAPTOR assigns dedicated hardware modules to each block, enabling more effective design and optimization. The full source code 111https://github.com/Kabir-Ehsan/Transformer_on_FPGA to reproduce the presented results or improve the design.
In summary, this work makes the following contributions:
@@
A novel accelerator architecture for a complete transformer that maximizes DSP and LUT utilization to enhance parallel processing and achieve low latency, supported by an analytical model for pre-execution estimates of resource use and latency.
@@
A modular design approach implemented using parameterized HLS codes to accommodate varying computation and data access patterns, as well as to allow design-time modification of different TNN components.
A runtime adaptive feature allows software-driven parameter adjustments to run different models without hardware re-synthesis.
2 Related Work
Various custom and partially adaptive FPGA accelerators have been developed for TNNs. Peng et al. [13, 27] introduced a coherent sequence length–adaptive algorithm-hardware co-design for Transformer acceleration and explored column-balanced block-wise pruning. Qi et al. [19, 34] proposed an acceleration framework combining balanced model compression at the algorithm level with hardware-level FPGA optimization. Chen et al. [35] developed an analytical model for evaluating spatial TNN accelerators, considering FPGA compute and memory resources, identifying optimal parallelization and buffering strategies, and providing reusable HLS kernels. Similarly, we developed an analytical model to estimate the latency and resource utilization of ADAPTOR and designed it modularly, with each module implemented as an HLS function for easy optimization and reuse. Qin et al. [36] designed a TNN accelerator with separate attention and linear kernels for long input sequences, applying tiling only to the attention layer, whereas our design applies unique tiling strategies to both attention and linear layers. Both architectures incorporate analytical models. The energy-efficient FTRANS framework [18] employed an improved block-circulant matrix method for algorithm-level sparsity, alongside a dedicated accelerator designed for this approach. Most of these architectures target specific TNNs and sparsity patterns, lacking runtime flexibility to reconfigure the computing structure for different applications. In contrast, ADAPTOR can be programmed from software for any dense TNN model. FlexRun [37] identified key NLP model components, implemented them on a state-of-the-art FPGA accelerator, performed design space exploration to determine the optimal architecture for a given NLP model, and enabled automatic reconfiguration based on the results. FET-OPU [38] presented an overlay architecture for general TNN acceleration featuring a DSP-packed Matrix Multiplication Unit (MMU) with a FIFO-based data caching mechanism. FlightLLM [16] introduced a configurable sparse DSP chain for handling diverse sparsity patterns efficiently, an always-on-chip decode scheme for improved memory bandwidth with mixed-precision support, and a length-adaptive compilation method to minimize instruction storage overhead for large language models. TRAC [39] focused on dedicated hardware generation, integrating code generation into the compilation process to create parameterized and synthesized modules for specific Transformer configurations-unlike fixed, though parameterizable, overlays. EFA-Trans [31] supports both dense and sparse computation patterns but requires hardware resynthesis to switch between them. Moreover, none of these works examined optimal tile sizes or DSP utilization for maximum parallelism as done in ADAPTOR.
@@ 4 ADAPTOR’s Architecture
4.1 Attention Module
The overall architecture designed to accelerate the attention mechanism is illustrated in Fig. 2. It consists of three principal processing modules (PMs), denoted as , , and , according to the specific operations they perform. Each of these modules begins operation only after the previous module has completed its computations. This strict sequential execution ensures that all data dependencies are respected and simplifies control logic. The number of module instances corresponds to the number of attention heads (). Within each module, computation is carried out by an array of processing elements (PEs), where each PE incorporates a DSP48 unit responsible for multiplication and accumulation (MAC) operations. The organization of the PE arrays varies across modules, as their computational demands and data access patterns differ. To accommodate these differences, the modules are implemented as separate functions in high-level synthesis (HLS), thereby enabling targeted optimization of the corresponding register-transfer level (RTL) components. Parallel data access is supported by distributing input activations and weights across multiple BRAMs and LUTRAMs.
Each PE operates independently, equipped with its own local memory, control logic, and computational resources. The weight matrices associated with the generation of queries (), keys (), and values () are stored as two-dimensional arrays of dimension , where denotes the tile size of the attention module. This tiling strategy partitions the larger weight matrices into sub-matrices, thereby facilitating efficient parallelization. The interplay between the number of heads, tiling parameters, and the HLS array partitioning directives determines how these arrays are mapped onto multiple two-port BRAMs. Since BRAM ports are limited, careful partitioning and scheduling of data transfers ensure that all operands required concurrently by the DSP units are accessible without contention. The intermediate , , and matrices, each of size where denotes the sequence length, are buffered locally to support subsequent stages of computation.
4.1.1 QKVPM module:
The module is responsible for generating the query, key, and value matrices. It incorporates dedicated BRAMs for the weights (, , ) and for the input activations (), which provide parallel data access to the DSP units within the processing element (PE) array. To accommodate on-chip memory constraints, the weight and input arrays are divided into subarrays using a tiling strategy, ensuring that the data can be efficiently mapped onto BRAMs or LUTRAMs. The number of times the module is invoked is determined by the tiling factor, resulting in a total of iterations. At each iteration, the buffers for , , , and are populated with distinct tiles of data, after which computation is initiated within the PEs.
During these operations, the corresponding bias terms for the , , and matrices are fetched from off-chip memory into registers in parallel with the primary computations of the module. These biases are subsequently integrated into the generated matrices, thereby completing the linear transformations. The computational flow of this module is summarized in Algorithm 9 of Section 10, where pipelining of the outer loop facilitates full unrolling of the innermost loop. This design yields an array of PEs, thereby maximizing throughput while maintaining an efficient mapping of resources.
4.1.2 QKPM module:
The module carries out the matrix–matrix multiplication between the and matrices. Since these matrices are relatively small in dimension, tiling is not required. The computational flow is summarized in Algorithm 11 of Section 10, where full unrolling of the innermost loop produces processing elements (PEs). Within this module, the and matrices are buffered to enable parallel access by the DSP units. In addition to the multiplication operations, the division specified in Equation 1 is also performed within this module using LUT resources. To avoid excessive LUT utilization, the degree of parallelism for this operation is deliberately constrained. The output of this module is the intermediate attention weight matrix , which is stored in either BRAMs or registers depending on availability and access requirements. These weights are subsequently passed to the non-linear softmax function, implemented in HLS using LUTs and flip-flops, to complete the attention score computation.
4.1.3 SVPM module:
The normalized attention weight matrix (), obtained from the softmax operation, is supplied to the module, where it is combined with the value () matrix through matrix–matrix multiplication. As described in Algorithm 12 of Section 10, the innermost loop is fully unrolled, enabling processing elements to operate in parallel. The resulting output, referred to as the attention score, represents a weighted aggregation of the value vectors and constitutes the final contribution of the attention mechanism to the subsequent layers.
4.2 Feedforward Network Module
The architecture developed to accelerate the feedforward network (FFN) module is depicted in Fig. 3. Three distinct RTL modules , , and are implemented to support variations of the FFN across different architectural configurations. They are executed sequentially to respect the inherent data dependencies. Each module begins processing only after the preceding module has fully completed its computation. In high-level synthesis (HLS), these modules are described as separate functions, each defined by input and output arrays of different dimensions, which are subsequently mapped onto BRAMs or LUTRAMs during synthesis. Since the computational workload differs across the modules, each function is optimized independently, resulting in varying numbers of processing elements depending on the unrolling factor applied to the innermost loop. The weights of the FFN are stored in a two-dimensional array () of dimensions , where denotes the tile size in FFN. This tiling strategy partitions the weight matrices into smaller blocks, facilitating parallel access and efficient memory utilization. Among the three RTL modules, both and are followed by layer normalization (LN), ensuring stabilized activations before passing results to subsequent stages.
4.2.1 PM module
The module performs the initial linear transformation on the attention scores, serving as the first stage of the feedforward network. To accommodate on-chip memory constraints, the arrays used by the processing elements (PEs) are tiled along both dimensions. Consequently, the module is invoked times to complete the transformation. As outlined in Algorithm 14 of Section 10, pipelining of the second loop enables full unrolling of the innermost loop (line 7), producing PEs in total. This corresponds to parallel computational units.
4.2.2 PM module
Building upon the normalized outputs of , the module performs the second linear transformation, expanding the intermediate representation. Similar to , arrays are tiled along both dimensions, though this module requires accesses due to the increased dimensionality of the operation. The computational flow is summarized in Algorithm 15 of Section 10, where pipelining again enables full unrolling of the innermost loop (line 7). This results in PEs, corresponding to units of parallelism, consistent with the structural design of the first module.
4.2.3 PM module
The module applies the final linear transformation to the normalized outputs of , projecting them back to the original model dimension. As with the preceding modules, arrays are tiled along both dimensions, requiring iterations to complete the computation. Algorithm 10 of Section 10 describes the workflow, where pipelining and full loop unrolling (line 7) yield PEs. This corresponds to , reflecting the higher dimensionality of this stage.
@@ 4.3 Load Weights Unit
Three dedicated Load_Weights units are employed to manage the transfer of parameters from external memory to on-chip buffers. The first unit supplies weights to the weight memories of the attention heads (Fig. 2), while the second serves the feedforward network (Fig. 3). A third unit is responsible for loading the weights associated with the layer normalization modules. This separation ensures that weight data can be delivered efficiently to each functional block in accordance with its computational demands. For the attention module, weights are represented in HLS as two-dimensional arrays of dimension , where denotes the tile size applied to partition the larger matrices into sub-matrices. After synthesis, these arrays are mapped onto dual-port BRAMs or LUTRAMs, and are populated iteratively with tile-specific data transferred from external memory at each iteration. In the feedforward network, the weight arrays are defined with dimensions . Here, corresponds to the tiling parameter, defined as , while equals . These weights are therefore partitioned along both row and column dimensions, requiring iterative loading across tiles. As in the attention module, they are synthesized as dual-port BRAMs or LUTRAMs. The weights for layer normalization are comparatively simple, represented as one-dimensional arrays of length . As no tiling is required in this case, the entire weight set is transferred in a single step and subsequently synthesized into dual-port BRAMs or LUTRAMs. The complete procedure for weight loading is outlined in Algorithm 1 of Section 10.
@@ 4.4 Load Inputs Unit
Input data are transferred from external memory into dedicated input BRAMs, which are implemented in HLS as dual-port, two-dimensional arrays of size , where denotes the sequence length. These BRAMs are reused across encoder and decoder layers, allowing each layer to access the outputs of the previous layer as inputs for subsequent computations. Three distinct Load_inputs units manage the data movement to accommodate differences in computation, tiling, and array dimensions across modules. The first unit populates the intermediate input BRAMs of each attention head (Fig. 2) using Algorithm 2 of Section 10. These BRAMs are represented as two-dimensional arrays of size , where tiling is applied along the column dimension. Consequently, data are loaded iteratively times to supply all columns to the processing elements.
The second unit transfers data to the Score BRAMs of the module (Fig. 3), defined as two-dimensional arrays of size , using Algorithm 3. The third unit supplies data to the LN1 buffers of the module (Fig. 3), represented as arrays of size and loaded according to Algorithm 4 of Section 10. The separation of load units ensures efficient handling of the differing computational demands, tile sizes, and array shapes across the attention and feedforward network modules.
@@ 4.5 Load Biases Unit
Bias parameters are stored in registers due to their relatively small size, enabling low-latency access during computation. Three dedicated Load_bias units manage the transfer of biases from external memory to the corresponding registers. The first unit supplies biases to the registers of each attention head in accordance with Algorithm 5. The same procedure, as described in Algorithm 6 of Section 10, is used to load biases for the feedforward network and the layer normalization modules. In HLS, biases are represented as one-dimensional arrays, and the application of a complete array partition pragma maps these arrays directly to registers. Since tiling is unnecessary for these small vectors, each array is loaded in a single transfer, providing all bias values simultaneously.
@@ 4.6 Activation Unit
The activation functions employed within the transformer architecture are implemented at this stage. Commonly used functions include ReLU, GeLU, and softmax, each defined according to its mathematical formulation. After synthesis, these functions are realized using LUTs to support efficient hardware computation. While the implementations of ReLU and GeLU are straightforward, the softmax function involves more complex operations; therefore, only its implementation is detailed in Algorithm 7 of Section 10.
@@ 4.7 Layer Normalization Unit
The Layer Normalization (LN) unit computes the mean and variance of the outputs from both the attention and feedforward network layers, following Equation 4. These statistics are then used to normalize the outputs, which are subsequently scaled by learned weights and shifted by biases in an element-wise manner. Notably, the outputs of and are processed through the LN unit prior to subsequent stages, as illustrated in Fig. 3. The corresponding HLS implementation is provided in Algorithm 8 of Section 10.